OpenStudy (anonymous):

walk me through how 2 solve this please

OpenStudy (anonymous):

\[\frac{ 8 }{ x }+\frac{ 15 }{ x ^{2} }=-1\]

OpenStudy (anonymous):

multiply everithing by x^2

OpenStudy (anonymous):

every term by x^2

OpenStudy (anonymous):

the numerators 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

@ajprincess ............

OpenStudy (ajprincess):

\[x^2*\frac{8}{x}+x^2*\frac{15}{x^2}=-1*x^2\]

OpenStudy (anonymous):

what do u do after that?

OpenStudy (ajprincess):

\[x*\cancel x*\frac{8}{\cancel x}+\cancel x^2*\frac{15}{x^2}=-x^2\] \[8x+15=-x^2\]

OpenStudy (ajprincess):

getting this?

OpenStudy (ajprincess):

\(8x+15=-x^2\) \(8x+15+x^2=0\) \(x^2+8x+15=0\) Nw let us use the quadratic formula to find x. \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (ajprincess):

Compare the equation \(x^2+8x+15=0\) with \(ax^2+bx+c=0\) nd cam u tell me the values of a, b and c?

OpenStudy (anonymous):

yeah a=0 b=8 c=15

OpenStudy (ajprincess):

b nd c r right. pls check a.

OpenStudy (anonymous):

1?

OpenStudy (ajprincess):

yup:)

OpenStudy (ajprincess):

Nw can u find x?

OpenStudy (anonymous):

ill try, 1 sec

OpenStudy (ajprincess):

ok. take ur time:)

OpenStudy (anonymous):

-8?

OpenStudy (ajprincess):

nope:(

OpenStudy (ajprincess):

\[x=\frac{-8\pm\sqrt{8^2-(4*1*15)}}{2*1}\]

OpenStudy (anonymous):



OpenStudy (ajprincess):

if u dnt mind can u pls make it a bit clear?

OpenStudy (anonymous):

untill the last part i was simplifying the sq root. i got 2 and then added the -8 +- and the devide by 2, making it -8(-+) 2/2 i thought the 2's cancle each other out making it 8

OpenStudy (anonymous):

*-8

OpenStudy (ajprincess):

nope. this is what u get finally right \[x=\frac{-8\pm2}{2}\] that is \(\Large{ x=\frac{-8+2}{2}}\) or \(\Large{ x=\frac{-8-2}{2}}\) first u need to simplify the numerator.

OpenStudy (anonymous):

ok -6 or -10?

OpenStudy (ajprincess):

those are ur numerators right?

OpenStudy (anonymous):

i think so.........

OpenStudy (ajprincess):

\[x=\frac{-8+2}{2}=\frac{-6}{2}=?\]

OpenStudy (anonymous):

-6/2= -3 -5?

OpenStudy (ajprincess):

-6/2=-3? or -5?

OpenStudy (ajprincess):

@janeDoe100

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

all of the answers on my paper have 3 and 5 in it. how?

OpenStudy (ajprincess):

-3 is right:) Nw tell me what is \[x=\frac{-8-2}{2}=\frac{-10}{2}=?\]

OpenStudy (anonymous):

-5

OpenStudy (ajprincess):

ya right:) so ur x values are -3 and -5. tell me the options given to u?

OpenStudy (anonymous):

-3 and -5 is b :D

OpenStudy (ajprincess):

:)